### v2

@@ I Introduction
One of the most striking properties in topological phases of matter is the appearance of non-trivial edge states Halperin (1982); Wen (1990); Stone (1991); Frohlich and Kerler (1991). The most well-known example is the Integer Quantum Hall (IQH) effect, where non-trivial bulk Chern number leads to the presence of gapless edge state with U(1) symmetry Hatsugai (1993), which can be probed by the Hall conductance $\sigma_{H}$. Even in the absence of U(1) symmetry, systems with nonzero chiral central charge $c_{-}$, which signals nonzero thermal Hall conductivity, have a gapless edge state Kane and Fisher (1997). The thermal or electric Hall conductivity can be characterized through the bulk effective theory, as well as extracted from a microscopic wavefunction, through various entanglement measures such as modular commutator, through randomized measurement, and so on Kitaev (2006); Mitchell et al. (2018); Tu et al. (2013); Zaletel et al. (2013); Kim et al. (2022a, b); Zou et al. (2022); Fan (2022); Fan et al. (2022); Shiozaki et al. (2018); Dehghani et al. (2021); Cian et al. (2021).
@@
While both the higher central charge and higher Hall conductivity are characterized through effective topological field theory, it is an important question to ask if these quantities allow microscopic characterizations. In the prior work by the authors Kobayashi et al. (2024a), it was found that the higher central charge can be extracted from a single wavefunction, by evaluating the expectation value of a unitary called the partial rotation. The extracted higher central charge then allows us to constrain the low-energy spectrum of the bulk-boundary system, e.g., completely determine if the given wavefunction bosonic of Abelian topological order has a gappable edge state or not. In this work, we study microscopic characterization of the higher Hall conductivity in the presence of $\mathrm{U}(1)^{f}$ symmetry, and explore its physical consequences including constraints on the low energy spectrum of edge states enforced by symmetry.
@@ I.1 Summary of results
Let us consider a single wave function of a fermionic FQH state on a cylinder |$\ket{\Psi_{a}}$ with a fixed topological sector, which is labeled by a simple anyon of the bulk topological order $a$. We then take the expectation value of the “partial rotation” unitary, which is the product of the partial spatial rotation $T_{\theta,\mathrm{A}}$ associated with the partial $\mathrm{U}(1)$ transformation $U_{\phi,\mathrm{A}}$ on the wave function,
@@
In a slight abuse of notation, we also refer to the expectation values as partial rotation when it is clear from context. The case of $\phi=0$ reduces to the partial rotation operator introduced in Ref. Kobayashi et al., 2024a. We show that a nontrivial phase of the above quantity $\mathcal{T}_{a}\left(\theta,\phi\right)$ with specific rotation angle is an obstruction to having a gapped edge state while preserving $\mathrm{U}(1)^{f}$ symmetry. That is, for integers $n$ satisfying the condition $\gcd(n,{N_{\mathrm{FS}}})=1$ with ${N_{\mathrm{FS}}}$ an integer called the Frobenius-Schur exponent, we demonstrate that
$\mathcal{T}_{1}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ $\propto\exp[\frac{2\pi i}{n}\bigg{(}\frac{\sigma_{H}}{8}-\Big{(}\frac{n^{2}+2}{24}\Big{)}c_{-}\bigg{)}\biggr{]}\zeta_{n}^{H}.$ (3)
This quantity is the combination of $\sigma_{H}$, $c_{-}$, and $\zeta_{n}^{H}$. The latter gives an obstruction to the presence of a symmetric gapped edge state beyond $\sigma_{H}$, $c_{-}$.
@@
Next, let us describe our result valid for Abelian topological order. We will first see that both electric and thermal Hall conductivity can be extracted by partial rotation and topological entanglement entropy of a single wavefunction. This is valid in either fermionic or bosonic Abelian phases, and we numerically verify the result on a bosonic Laughlin states. Even when $\sigma_{H}=c_{-}=0$, we show that once the topological entanglement entropy of the given wave function is known, the partial rotations completely determine if a given wave function for Abelian topological order admits a gapped edge state preserving $\mathrm{U}(1)$ symmetry or not. This statement is again valid for either bosonic or fermionic Abelian phases.
@@ II.1 Higher central charge of bosonic topological order
The higher central charge $\zeta_{n}$ is labeled by a positive integer $n$. When $n$ satisfies $\gcd(n,{N_{\mathrm{FS}}})=1$, higher central charge is defined as
@@
The trivial higher central charge $\zeta_{n}=1$ for all $n$ such that $\gcd(n,{N_{\mathrm{FS}}})=1$ gives necessary conditions to gapped boundary of bosonic topological order, which are generally independent of the vanishing chiral central charge condition $c_{-}=0$. This implies that the set of quantities $\{\zeta_{n}\}$ prohibits the existence of the gapped edge state even if $c_{-}$ is vanishing. In the special case with $n=1$, $\zeta_{1}$ just gives the Gauss-Milgram formula for $c_{-}$ mod 8; $\zeta_{1}=e^{2\pi ic_{-}/8}$. In this sense, $\zeta_{n}$ with $n>1$ can naturally be regarded as a “higher” version of $c_{-}$, hence the name of higher central charge.
When the topological order is Abelian, $\zeta_{n}=1$ for all $n$ satisfying $\gcd\bigl{(}n,\frac{{N_{\mathrm{FS}}}}{\gcd(n,{N_{\mathrm{FS}}})}\bigr{)}=1$ gives necessary and sufficient conditions for the gapped boundary Kaidi et al. (2021), so $\{\zeta_{n}\}$ completely determines the gappability of the edge state of Abelian topological order.  [^1] Though we defined $\zeta_{n}$ only when $\gcd(n,{N_{\mathrm{FS}}})=1$ in Eq. (4), we can define $\zeta_{n}$ by the same formula as (4) for $n$ satisfying $\gcd(n,\frac{{N_{\mathrm{FS}}}}{\gcd(n,{N_{\mathrm{FS}}})})=1$ in the Abelian case.
@@ II.2 Higher Hall conductivity of fermionic FQH states
The higher Hall conductivity $\zeta_{n}^{H}$ is again labeled by positive integers $n$ satisfying $\gcd(n,{N_{\mathrm{FS}}})=1$. We define $\zeta_{n}^{H}$ as the following quantity computed from the data of anyons:
@@
Importantly, $\zeta_{n}^{H}=1$ for all $n$ gives necessary conditions to having a symmetry-preserving gapped edge state, generally independent of the most fundamental conditions $c_{-}=\sigma_{H}=0$. This implies that the set of quantities $\{\zeta^{H}_{n}\}$ prohibits the existence of the symmetry-preserving gapped edge states even if both $c_{-}$ and $\sigma_{H}$ are vanishing. In that sense, $\zeta^{H}_{n}$ is regarded as a “higher” version of thermal and electric Hall conductivity, hence the name of higher Hall conductivity.
@@
The higher Hall conductivity $\zeta_{n}^{H}$ cannot generally be expressed as a linear combination of $c_{-},\sigma_{H}$. Meanwhile, in the special case with $n=1$, we have the relation
@@ III.1 Partial rotation for the fermionic phase
see Fig. 1 for a schematic illustration.
@@
We note the viability of measuring the partial rotation. If the wavefunction is put on a universal quantum computer, we can evaluate the expectation value of the unitary operator through the Hadamard test.
@@ III.1.1 Extraction of Hall conductivity via partial rotation in fermionic Abelian phases
When the phase is Abelian, the Frobenius-Schur exponent ${N_{\mathrm{FS}}}$ must divide $2\mathcal{D}^{2}$. By using this fact and $\zeta_{1}^{H}=e^{\frac{2\pi i}{8}(c_{-}-\sigma_{H})}$, we can obtain $c_{-}$ and $\sigma_{H}$ from the following two quantities
@@ III.2 Bosonic phase
When $m=0$ and $a=1$, it reduces to the higher central charge of the bosonic topological phase Kobayashi et al. (2024a)
@@ IV Numerical results
We first briefly describe the procedure to compute the partial $\mathrm{U}(1)$ rotation given a matrix product state (MPS). The technology we need here is very similar to the one used in Ref. Kobayashi et al., 2024a, whereas the only difference is that we also need the particle number label for the $\mathrm{U}(1)$ rotation. For the quantum Hall systems we consider here, we always work with an infinite cylindrical geometry, and Landau gauge orbitals around the cylinder. We conserve both momentum around the cylinder and particle number, allowing us to reach bond dimension $\chi=3200$ and cylinder circumference $L_{y}=40\ell_{B}$ such that the partial rotation in all Abelian phases studied is well converged. We use the real-space entanglement spectrum (RSES) algorithm described in Ref. Zaletel and Mong, 2012 to cut the cylinder in two halves in physical space, along which we perform the partial $\mathrm{U}(1)$ rotation. The final expectation value we need can be evaluated at the virtual bond corresponding to the cut,
@@
The non-Abelian phases, however, present more numerical challenges. Here we show $\mathcal{T}_{1}^{f}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ of the Moore-Read state in Fig. 3. The effective field theory for the Moore-Read state is obtained by condensing a fermion $\psi$ in the $\mathrm{U}(1)_{8}\times\mathrm{Ising}$ topological order. The anyons are labeled by $([j]_{8},a)$ with $a\in\{1,\sigma,\epsilon\}$, where the fermion is labeled by $\psi=([4]_{8},\epsilon)$. The fractional U(1) charge is given by $Q_{([j]_{8},a)}=j/4$. See Table 3 for spins of the anyons in the Moore-Read state, and Table 1 for the values of $\mathcal{T}_{1}^{f}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ for small $n$. In the expression of $\mathcal{T}_{1}^{f}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ for even $n$ in Eq. (12), we need to sum over the ${\mathbb{Z}}_{2}$ vortices $v^{\prime}$ with the smallest spin, which are $v^{\prime}=([1]_{8},1),([7]_{8},1),([0]_{8},\sigma)$ with $h_{m}=1/16$ in $\mathrm{U}(1)_{8}\times\mathrm{Ising}$.
@@
Finally, we show results of bosonic Abelian phases and explore the possibility to extract electric Hall conductance from partial rotation. Here we present $\mathcal{T}_{1}^{b}\left(\frac{2\pi}{n},\frac{2\pi}{n}\right)$ of the bosonic $\nu=1/2$ Laughlin state, with a single non-trivial anyon being the semion. We follow Ref. Kobayashi et al., 2024a to obtain the ground state MPS by solving the half-filled LLL of two-dimensional bosons with a contact interaction $V_{0}=1$ plus a small perturbation $\delta V_{2}=0.1$ with iDMRG. As shown in Fig. 4 (a), $\mathcal{T}_{1}^{b}\left(\frac{2\pi}{n},\frac{2\pi}{n}\right)$ always converges to the expected phase as shown in Table 2 at sufficiently large $L_{y}$. Combining with $\mathcal{T}_{1}^{b}\left(\frac{2\pi}{n},0\right)$ obtained in Ref. Kobayashi et al., 2024a, we can extract the electric Hall conductance using Eq. (21). As shown in Fig. 4 (b), the Hall conductance converges to the expected value $\sigma_{H}=1/2$. The numerical success of Eq. (21) presents partial rotation as an appealing method to extract boson Hall conductance.

### v3

@@ I Introduction
One of the most striking properties in topological phases of matter is the appearance of non-trivial edge states Halperin (1982); Wen (1990); Stone (1991); Frohlich and Kerler (1991). The most well-known example is the Integer Quantum Hall (IQH) effect, where non-trivial bulk Chern number leads to the presence of gapless edge state with U(1) symmetry Hatsugai (1993), which can be probed by the Hall conductance $\sigma_{H}$. Even in the absence of U(1) symmetry, systems with nonzero chiral central charge $c_{-}$, which signals nonzero thermal Hall conductivity, have a gapless edge state Kane and Fisher (1997). The invariants including thermal or electric Hall conductivity can be characterized through the bulk effective theory, as well as extracted from a microscopic wavefunction, through various entanglement measures such as modular commutator, through randomized measurement, and so on Kitaev (2006); Mitchell et al. (2018); Tu et al. (2013); Zaletel et al. (2013); Kim et al. (2022a, b); Zou et al. (2022); Fan (2022); Fan et al. (2022); Shiozaki et al. (2018); Dehghani et al. (2021); Cian et al. (2021); Liang et al. (2024a, b); Kobayashi et al. (2024a).
@@
While both the higher central charge and higher Hall conductivity are characterized through effective topological field theory, it is an important question to ask if these quantities allow microscopic characterizations. In the prior work by the authors Kobayashi et al. (2024b), it was found that the higher central charge can be extracted from a single wavefunction, by evaluating the expectation value of a unitary called the partial rotation. The extracted higher central charge then allows us to constrain the low-energy spectrum of the bulk-boundary system, e.g., completely determine if the given wavefunction bosonic of Abelian topological order has a gappable edge state or not. In this work, we study microscopic characterization of the higher Hall conductivity in the presence of $\mathrm{U}(1)^{f}$ symmetry, and explore its physical consequences including constraints on the low energy spectrum of edge states enforced by symmetry.
@@ I.1 Summary of results
Let us consider a single wave function of a fermionic FQH state on a cylinder |$\ket{\Psi_{a}}$ with a fixed topological sector, which is labeled by a simple anyon of the bulk topological order $a$. We take the bipartition of the cylinder into the subsystem A and B, see Fig. 1. We then compute the expectation value of the “partial rotation” unitary, which is the product of the partial spatial rotation $T_{\theta,\mathrm{A}}$ associated with the partial $\mathrm{U}(1)$ transformation $U_{\phi,\mathrm{A}}$ on the wave function,
@@
In a slight abuse of notation, we also refer to the expectation values as partial rotation when it is clear from context. The case of $\phi=0$ reduces to the partial rotation operator introduced in Ref. Kobayashi et al., 2024b. We show that a nontrivial phase of the above quantity $\mathcal{T}_{a}\left(\theta,\phi\right)$ with specific rotation angle is an obstruction to having a gapped edge state while preserving $\mathrm{U}(1)^{f}$ symmetry. That is, for integers $n$ satisfying the condition $\gcd(n,{N_{\mathrm{FS}}})=1$ with ${N_{\mathrm{FS}}}$ an integer called the Frobenius-Schur exponent, we demonstrate that
$\mathcal{T}_{1}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ $\propto\exp[\frac{2\pi i}{n}\bigg{(}\frac{\sigma_{H}}{8}-\Big{(}\frac{n^{2}+2}{24}\Big{)}c_{-}\bigg{)}\biggr{]}\zeta_{n}^{H},$ (3)
where $\propto$ in this paper always means being proportional up to a positive real number. This quantity is the combination of $\sigma_{H}$, $c_{-}$, and $\zeta_{n}^{H}$. The latter gives an obstruction to the presence of a symmetric gapped edge state beyond $\sigma_{H}$, $c_{-}$.
@@
Next, let us describe our result valid for Abelian topological order. We will first see that both electric and thermal Hall conductivity can be extracted by partial rotation and topological entanglement entropy of a single wavefunction. [^1] Throughout the paper, we assume that topological entanglement entropy gives total quantum dimension $\mathcal{D}$ of the topological order, i.e., there is no spurious contributions as discussed in Cano et al. (2015); Zou and Haah (2016); Williamson et al. (2019); Kim et al. (2023); Levin (2024). This assumption is valid when the entanglement spectrum is given by the edge CFT, which is also used to calculate the partial rotations in this paper. This is valid in either fermionic or bosonic Abelian phases, and we numerically verify the result on a bosonic Laughlin states. Even when $\sigma_{H}=c_{-}=0$, we show that once the topological entanglement entropy of the given wave function is known, the partial rotations completely determine if a given wave function for Abelian topological order admits a gapped edge state preserving $\mathrm{U}(1)$ symmetry or not. This statement is again valid for either bosonic or fermionic Abelian phases.
@@ II.1 Higher central charge of bosonic topological order
The higher central charge $\zeta_{n}$ is labeled by a positive integer $n$, which is defined as
@@
The trivial higher central charge $\zeta_{n}=1$ for all $n$ such that $\gcd(n,{N_{\mathrm{FS}}})=1$ gives necessary conditions to gapped boundary of bosonic topological order, which are generally independent of the vanishing chiral central charge condition $c_{-}=0$ Ng et al. (2019). This implies that the set of quantities $\{\zeta_{n}\}$ prohibits the existence of the gapped edge state even if $c_{-}$ is vanishing. In the special case with $n=1$, $\zeta_{1}$ just gives the Gauss-Milgram formula for $c_{-}$ mod 8; $\zeta_{1}=e^{2\pi ic_{-}/8}$. In this sense, $\zeta_{n}$ with $n>1$ can naturally be regarded as a “higher” version of $c_{-}$, hence the name of higher central charge.
When the topological order is Abelian, $\zeta_{n}=1$ for all $n$ satisfying $\gcd\bigl{(}n,\frac{{N_{\mathrm{FS}}}}{\gcd(n,{N_{\mathrm{FS}}})}\bigr{)}=1$ gives necessary and sufficient conditions for the gapped boundary Kaidi et al. (2021), so $\{\zeta_{n}\}$ completely determines the gappability of the edge state of Abelian topological order.
@@ II.2 Higher Hall conductivity of fermionic FQH states
The higher Hall conductivity $\zeta_{n}^{H}$ is again labeled by positive integers $n$, which is defined as the following quantity computed from the data of anyons:
@@
Importantly, $\zeta_{n}^{H}=1$ for all $n$ satisfying $\mathrm{gcd}(n,{N_{\mathrm{FS}}})=1$ gives necessary conditions to having a symmetry-preserving gapped edge state, generally independent of the most fundamental conditions $c_{-}=\sigma_{H}=0$ Kobayashi (2022). This implies that the set of quantities $\{\zeta^{H}_{n}\}$ prohibits the existence of the symmetry-preserving gapped edge states even if both $c_{-}$ and $\sigma_{H}$ are vanishing. In that sense, $\zeta^{H}_{n}$ is regarded as a “higher” version of thermal and electric Hall conductivity, hence the name of higher Hall conductivity.
@@
The higher Hall conductivity $\zeta_{n}^{H}$ cannot generally be expressed as a linear combination of $c_{-},\sigma_{H}$. Meanwhile, in the special case with $n=1$, we have the relation Lapa and Levin (2019)
@@ III.1 Partial rotation for the fermionic phase
see Fig. 1 for a schematic illustration. The operator $T_{\theta,\mathrm{A}}$ is referred to as a partial rotation operator in this paper, while we note that it is called partial translation in previous work e.g., Tu et al. (2013).
@@
We note the viability of measuring the partial rotation. If the wavefunction is put on a universal quantum computer, we can evaluate the expectation value of the unitary operator through the Hadamard test as follows: Considering preparing the following state:
$HU^{\mathrm{cont}}\ket{+}_{\mathrm{ancilla}}\ket{\psi}$ (14)
where $H$ is the Hadarmard gate, $U^{\mathrm{cont}}$ is the controlled $U$ operation, |±$\ket{\pm}_{\mathrm{ancilla}}=(\ket{\uparrow}_{\mathrm{ancilla}}\pm\ket{\downarrow}_{\mathrm{ancilla}})/\sqrt{2}$ is an ancilla qubit, and |$\ket{\psi}$ is the state of interest. Writing out the action of $U^{\mathrm{cont}}$ explicitly, the state we get is
[…]
The $\sigma^{z}$ expectation value of the ancilla qubit is then given by $\mathrm{Re}(\innerproduct{\psi|U|\psi}{\psi|U|\psi})$, which is the real part of the expectation value. We can similarly measure the imaginary part by using (|$(\ket{\uparrow}_{\mathrm{ancilla}}-i\ket{\downarrow}_{\mathrm{ancilla}})/\sqrt{2}$ as the initial state on the ancilla register. By choosing $U$ to be the partial rotation operator, we can measure the expectation values straightforwardly on a quantum computer.
@@ III.1.1 Extraction of Hall conductivity via partial rotation in fermionic Abelian phases
When the phase is Abelian, the Frobenius-Schur exponent ${N_{\mathrm{FS}}}$ must divide $2\mathcal{D}^{2}$. By using this fact and $\zeta_{1}^{H}=e^{\frac{2\pi i}{8}(c_{-}-\sigma_{H})}$, we can obtain $c_{-}$ and $\sigma_{H}$ from the following two quantities [^2] We note that Eqs. (16), (17) do not fully determine the values of $c_{-},\sigma_{H}$ as rational numbers, since the partial rotations depend on these quantities through the phases. However, given that the phases in Eqs. (16), (17) depend on $c_{-},\sigma_{H}$ with the fractions of large denominators $\mathrm{lcm}(2\mathcal{D}^{2}-1,8)$, it is expected that these two equations fix $c_{-},\sigma_{H}$ modulo the integers as large as the denominators $\mathrm{lcm}(2\mathcal{D}^{2}-1,8)$. If we want to get $c_{-},\sigma_{H}$ with higher precisions, the partial rotations with rotation angles $2\pi/(2m\mathcal{D}^{2}\pm 1)$ with $m\in\mathbb{Z}$ to give the additional information to these quantities, and it is expected that we can get $c_{-},\sigma_{H}$ with arbitrarily high precisions. It would be interesting to rigorously verify these exceptions.
@@ III.2 Bosonic phase
When $m=0$ and $a=1$, it reduces to the higher central charge of the bosonic topological phase Kobayashi et al. (2024b)
@@ IV Numerical results
We first briefly describe the procedure to compute the partial $\mathrm{U}(1)$ rotation given a matrix product state (MPS). The technology we need here is very similar to the one used in Ref. Kobayashi et al., 2024b, whereas the only difference is that we also need the particle number label for the $\mathrm{U}(1)$ rotation. For the quantum Hall systems we consider here, we always work with an infinite cylindrical geometry, and Landau gauge orbitals around the cylinder. We conserve both momentum around the cylinder and particle number, allowing us to reach bond dimension $\chi=3200$ and cylinder circumference $L_{y}=40\ell_{B}$ such that the partial rotation in all Abelian phases studied is well converged. We use the real-space entanglement spectrum (RSES) algorithm described in Ref. Zaletel and Mong, 2012 to cut the cylinder in two halves in physical space, along which we perform the partial $\mathrm{U}(1)$ rotation. The final expectation value we need can be evaluated at the virtual bond corresponding to the cut,
@@
The non-Abelian phases, however, present more numerical challenges. Here we show $\mathcal{T}_{1}^{f}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ of the Moore-Read state in Fig. 3. The set of anyons in the Moore-Read state is obtained by condensing a fermion $\psi$ in the $\mathrm{U}(1)_{8}\times\mathrm{Ising}$ topological order. The anyons are labeled by $([j]_{8},a)$ with $a\in\{1,\sigma,\epsilon\}$, where the fermion is labeled by $\psi=([4]_{8},\epsilon)$. The fractional U(1) charge is given by $Q_{([j]_{8},a)}=j/4$. See Table 3 for spins of the anyons in the Moore-Read state, and Table 1 for the values of $\mathcal{T}_{1}^{f}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ for small $n$. In the expression of $\mathcal{T}_{1}^{f}\left(\frac{2\pi}{n},\frac{\pi}{n}\right)$ for even $n$ in Eq. (12), we need to sum over the ${\mathbb{Z}}_{2}$ vortices $v^{\prime}$ with the smallest spin, which are $v^{\prime}=([1]_{8},1),([7]_{8},1),([0]_{8},\sigma)$ with $h_{m}=1/16$ in $\mathrm{U}(1)_{8}\times\mathrm{Ising}$.
@@
Finally, we show results of bosonic Abelian phases and explore the possibility to extract electric Hall conductance from partial rotation. Here we present $\mathcal{T}_{1}^{b}\left(\frac{2\pi}{n},\frac{2\pi}{n}\right)$ of the bosonic $\nu=1/2$ Laughlin state, with a single non-trivial anyon being the semion. We follow Ref. Kobayashi et al., 2024b to obtain the ground state MPS by solving the half-filled LLL of two-dimensional bosons with a contact interaction $V_{0}=1$ plus a small perturbation $\delta V_{2}=0.1$ with iDMRG. As shown in Fig. 4 (a), $\mathcal{T}_{1}^{b}\left(\frac{2\pi}{n},\frac{2\pi}{n}\right)$ always converges to the expected phase as shown in Table 2 at sufficiently large $L_{y}$. Combining with $\mathcal{T}_{1}^{b}\left(\frac{2\pi}{n},0\right)$ obtained in Ref. Kobayashi et al., 2024b, we can extract the electric Hall conductance using Eq. (23). As shown in Fig. 4 (b), the Hall conductance converges to the expected value $\sigma_{H}=1/2$. The numerical success of Eq. (23) presents partial rotation as an appealing method to extract boson Hall conductance.